# High efficiency quantification of 90Sr contamination in cow milk after a nuclear accident

### v1

@@ Abstract
The minimum detectable activity (MDA) was estimated at 0.33 Bq/L under optimal conditions, demonstrating the protocol’s sensitivity for low-level detection. A comparative analysis with existing methods centrifugation-based approaches and Dowex resin techniques revealed that our protocol outperforms in both strontium recovery and organic matter elimination. Alternative methods showed lower recovery rates (68 2% for Guérin’s method, 65 6% for Dowex resin) and suffered from procedural drawbacks, such as incomplete organic matter removal.
Applying this methodology to compare samples from certified laboratories confirmed its robustness, with liquid scintillation spectroscopy radioactivity values doubling after 14 days, consistent with secular equilibrium between 90Sr and 90Y. While the protocol is optimized for milk, future research should explore its applicability to other food matrices. The high yield, reliability, and ease of implementation position this method as an effective tool for nuclear emergency response and routine radiological monitoring.
@@
Milk , Strontium-90 , Freeze-drying , Calcination , Chromatography , LSC[HEPIA]organization=Nuclear Physics & Chemistry Laboratories, HEPIA HES‐SO, University of Applied Sciences and Arts Western Switzerland,addressline=Rue de la Prairie 4, city=Geneva, postcode=1202, country=Switzerland
@@ 1 Introduction
In recent years, growing concerns over nuclear safety have intensified due to potential reactor accidents, unauthorized releases in the environment, and geopolitical conflicts involving nuclear weapons. In response, multiple European countries, alongside national and international agencies, have been developing environmental monitoring protocols for nuclear emergencies [1, 2]. Many of these protocols prioritize food safety, with milk being a key focus due to its widespread consumption and its critical role in children’s nutrition.
Radiocontaminated milk containing 90Sr presents significant challenges due to its potential to propagate contamination through the food chain [3]. 90Sr, a key radioisotope released during nuclear accidents, is a high-energy - emitter with calcium-mimicking chemical and biological properties. This similarity raises major health concerns, as 90Sr can be incorporated into teeth and bones, leading to long-term radiotoxic effects.
This work presents a safe, efficient and user-friendly protocol for quantifying 90Sr radioactivity in milk, following three key steps: elimination of organic matter, strontium separation and measurement of the extracted radioactivity. The protocol involves freeze-drying, followed by calcination at 700°C to remove organic content. Strontium is selectively separated using ion exchange chromatography and the extracted radioactivity is measured via liquid scintillation spectroscopy, with mass spectroscopy employed for yield quantification (see Fig. 1). Finally, the protocol’s accuracy and reliability are validated through comparisons with established methodologies from the literature.
@@ 2.1 Reagents and Apparatus
All the reagents used in this study (HCl, TCA, NaOH, Na2CO3, H2O2, HNO3, C2H2O4) as well as the natural Sr tracer (ROTI®Star, 1000 mg/L) used for yield measurements, were purchased from Carl Roth AG. Sr Resin® chromatography columns (2 mL) were obtained from Eichrom Technologies Inc, while the90Sr standard solution (5 mL, 1.85 MBq at 01.02.2022) was sourced from Eckert & Ziegler Nuclitec GmbH. UHT milk (3.5 fat) was used for this study.
@@ 2.2 Freeze drying
Freeze-drying or else lyophilization, or else, is the initial step of the developed protocol, designed to convert liquid milk into dry powder and prevent issues associated with liquid milk calcination. The freeze-dryer is equipped with a three-tiered support, allowing three large-diameter Pyrex crystallizers to be placed under a glass dome during the process. The wide surface area minimizes sample thickness, thereby reducing drying time. Prior to freeze-drying, the samples are pre-frozen in a freezer overnight (approximately 18 h). The drying progress can be monitored by checking for liquid residues or ice at the bottom of the crystallizer. Upon completion, the milk forms a dry, easily detachable block, yielding approximately 12 g of powdered milk from 100 mL of liquid milk. The percentage of evaporated matter over time for a 100 mL crystallizer is illustrated in the Fig. 2.
@@ 2.3 Calcination
The powdered milk was transferred into a porcelain beaker and heated in a muffle furnace overnight (approximately 18 hours). Initially, calcination was performed at 550°C. However, organic matter removal was incomplete, yielding grayish ash (see Fig. 3). The colour of the ash serves as a reliable indicator of residual organic materials, with bright white signifying complete organic matter removal. Furthermore, calcined ash obtained at 550°C clogged the resin column or significantly reduced the flow rate.
@@ 2.4 Ion Exchange chromatography
The process begins with preconditioning the column using the same reagent as the sample solution. Since the calcined ash is diluted in 10 mL of 8M HNO3, the column is preconditioned by adding 5 mL of 8M HNO3.The diluted sample is then passed through the column. A series of rinses follows to separate strontium (stable or radioactive) from most other elements, based on their retention behaviour (cf. Fig. 4), which varies with nitric acid concentration. The rinsing sequence begins with 15 mL of 8M HNO3, followed by 5 mL of 3M HNO3 mixed with 0.05M oxalic acid to eliminate potential actinides. This step is crucial, as oxalic acid facilitates the formation of soluble actinide complexes, preventing their retention on the Sr Resin® column [6]. Finally, 7 mL of 8M HNO3 is used to complete the rinsing process.
@@ 2.5 Liquid Scintillation Spectroscopy
Liquid scintillation measurements were performed using a Packard TriCarb 2900TR spectrometer. The radioactive liquid eluted from the chromatography column was mixed with Ultima Gold AB scintillation liquid inside a 20-mL borosilicate glass vial. Samples were measured immediately after the chromatography process (t=0) and again two weeks later (t=14 d) to ensure that the 90Sr/90Y pair reached secular equilibrium. This approach allowed for a clearer distinction of each radioisotope’s contribution to the final spectrum shape, thereby optimizing the final activity calculation (see Fig. 5).
@@ 2.6 Mass Spectrometry
The process consisted of two main steps. The first step involved calibrating the instrument using 5 aqueous solutions spiked with a stable strontium tracer at various concentrations. These samples were prepared and measured to generate a calibration curve for strontium concentrations ranging from 0.1 to 10 ppm (see Fig. 6).
@@
Moreover, ICP-MS enables the detection of trace amounts of other elements within the sample. This capability was used to measure the decontamination factor (see Table 2) and assess whether other elements could interfere with the activity values obtained from liquid scintillation spectroscopy. The decontamination factor was determined by comparing element concentrations before and after chromatography.
Element | Decontamination factor |
Na | 617 ± 99 |
Mg | 2721 ± 668 |
K | 1369 ± 232 |
Ca | 16 ± 4 |
Ba | 46 ± 2 |
Pb | 159 ± 59 |
Th | 25 ± 8 |
Cl | 46 ± 5 |
Zr | 10 ± 1 |
Mo | 106 ± 20 |
@@ 3 Results and Discussion
As part of an intercomparison measurement between several certified Swiss nuclear laboratories, a milk sample containing 90Sr at an unknown concentration below the release limit (1 Bq/mL according to ORaP [9]) was received. To replicate a similar 90Sr concentration to that expected in the received samples, reference milk samples were prepared with a concentration of 0.45 Bq/mL. The protocol was then applied, achieving a high efficiency of 1002% as measured by ICP-MS, demonstrating the method’s reproducibility and robustness. These results confirm that the developed protocol retains 100% of the added 90Sr, enabling accurate quantification through liquid scintillation spectroscopy.
@@
To further validate the protocol’s accuracy, a second LSC measurement was conducted at 20 and 30 days to track the increase in 90Y activity until secular equilibrium was reached. At this point, the absolute activity of all measured samples doubled and then stabilized, confirming the effective separation of strontium and yttrium during chromatography (see Fig. 7).
@@
The detection limit of 0.1 Bq was determined by measuring a series of 90Sr-spiked samples ranging from 0.1 to 1 Bq (see Fig. 8) using the liquid scintillation spectrometer.
@@ 4 Conclusions and Outlook
This study presents a robust and efficient protocol for the quantification of 90Sr in milk, addressing critical food safety concerns related to potential radioactive contamination. The proposed method integrates freeze-drying, high temperature calcination (at 700°C), ion exchange chromatography, and liquid scintillation spectroscopy, achieving a high recovery efficiency of 1002%, as determined by ICP-MS measurements. Theoretical calculations indicate that the protocol achieves a Minimum Detectable Activity (MDA) of 0.33 Bq/L under optimal conditions. Additionally, the protocol was experimentally validated at a concentration of 0.45 Bq/mL, demonstrating its practical suitability for detecting low-level 90Sr contamination in milk, particularly during nuclear emergencies.
Comparative testing with alternative methodologies, including protocols centrifugation-based and Dowex® resin-based protocols, highlights the superior performance of the developed method in terms of strontium recovery and organic matter removal. For example, the Guérin protocol yielded 682% efficiency, while the Dowex® resin method achieved 656%, both exhibiting limitations in either strontium retention or organic matter elimination (see Table 3). These findings confirm the robustness, efficiency, and practicality of the proposed method for routine radiological monitoring and nuclear emergency response applications.
Despite its high performance, the protocol has certain limitations. While calcination at 700°C ensures near-complete organic matter removal, the process is time-intensive and may limit throughput in large-scale emergency scenarios. Future studies could explore alternative calcination techniques or optimized reagent formulations to accelerate the process while maintaining sample integrity. Additionally, although this study focused on milk as a representative food matrix, further research is required to adapt the protocol for other food products, including vegetables, grains, and processed foods.
Advancements in radioactivity detection technologies, such as the integration of ultra-low-level scintillation counters, could further enhance detection sensitivity and expand the protocol’s applicability to very low contamination levels. These innovations, combined with the protocol’s demonstrated efficiency and reliability, establish a robust framework for monitoring radioactive contamination in food supplies. This study contributes to international nuclear safety standards and provides a scientifically validated tool for safeguarding public health during nuclear incidents.
@@ Acknowledgements
The authors would like to express their sincere gratitude to the inTECH institute of HEPIA for their invaluable support throughout this research. The provision of equipment and laboratory time to our nuclear physics and nuclear chemistry laboratories was instrumental in the successful completion of this study.

### v2

@@ Abstract
The minimum detectable activity (MDA) was estimated at 0.33 Bq/L under optimal conditions, demonstrating the protocol’s sensitivity for low-level detection. A comparative analysis with existing methods, centrifugation-based approaches and Dowex resin techniques revealed that our protocol outperforms in both strontium recovery and organic matter elimination. Alternative methods showed lower recovery rates (68 2% for Guérin’s method, 65 6% for Dowex resin) and suffered from procedural drawbacks, such as incomplete organic matter removal.
Applying this methodology to compare samples from certified laboratories confirmed its robustness, with liquid scintillation spectroscopy radioactivity values doubling after 14 days, consistent with secular equilibrium between 90Sr and 90Y. While the protocol is optimized for milk, future research should explore its applicability to other food matrices. The high yield, reliability, and ease of implementation, position this method as an effective tool for nuclear emergency response and routine radiological monitoring.
@@
Milk , Strontium-90 , Freeze-drying , Calcination , Chromatography , LSC[HEPIA]organization=Nuclear Physics & Chemistry Laboratories, HEPIA HES‐SO, University of Applied Sciences and Arts Western Switzerlanda,addressline=Rue de la Prairie 4, city=Geneva, postcode=1202, country=Switzerland
@@ 1 Introduction
In recent years, growing concerns over nuclear safety have intensified due to potential reactor accidents, unauthorized releases in the environment, and geopolitical conflicts involving nuclear weapons. In response, multiple European countries, alongside national and international agencies, have been developing environmental monitoring protocols for nuclear emergencies [1, 2]. Many of these protocols prioritize food safety, with a particular emphasis on milk, given its widespread consumption and its essential role in child nutrition.
Radiocontaminated milk containing 90Sr presents significant challenges due to its potential to propagate contamination throughout the food chain [3]. 90Sr, a major radioisotope released during nuclear accidents, is a high-energy - emitter with calcium-mimicking chemical and biological properties. This mimicry raises major health concerns, as 90Sr can become incorporated into teeth and bone tissue, leading to long-term radiotoxic effects.
This work presents a safe, efficient and user-friendly protocol for quantifying 90Sr radioactivity in milk, structured around three key steps: elimination of organic matter, selective separation of strontium and measurement of the extracted radioactivity. The protocol begins with freeze-drying, followed by calcination at 700°C to remove organic content. Strontium is then selectively isolated using ion exchange chromatography and the radioactivity is measured via liquid scintillation spectroscopy, with mass spectroscopy employed for yield quantification (see Fig. 1). The protocol’s accuracy and reliability are validated through comparisons with established methods reported in the literature.
@@ 2.1 Reagents and Apparatus
All the reagents used in this study (HCl, TCA, NaOH, Na2CO3, H2O2, HNO3, C2H2O4) as well as the natural Sr tracer (ROTI®Star, 1000 mg/L) used for yield measurements, were purchased from Carl Roth AG. Sr Resin® chromatography columns (2 mL) were obtained from Eichrom Technologies Inc, while the90Sr standard solution (5 mL, 1.85 MBq at 01.02.2022) was sourced from Eckert & Ziegler Nuclitec GmbH. UHT full-fat cow milk (3.5 fat) was selected for this study to evaluate the protocol under conservative and challenging conditions, as the higher fat content increases the complexity of sample preparation. In contrast, using skimmed or semi-skimmed milk would likely simplify both calcination and purification steps, due to lower organic load.
@@ 2.2 Freeze drying
Freeze-drying or else lyophilization, is the initial step of the developed protocol, used to convert liquid milk into dry powder and prevent issues associated with liquid milk calcination. The freeze-dryer used in this study is equipped with a three-tiered support system, allowing simultaneous processing of three large-diameter Pyrex crystallizers placed under a glass dome. The wide surface area minimizes sample thickness, thereby reducing drying time. Prior to freeze-drying, the milk samples are pre-frozen in a freezer overnight (approximately 18 hours). The drying progress is monitored by visually inspecting the bottom of the crystallizer for the presence of residual liquid or ice. Upon completion, the milk forms a dry, easily detachable block, yielding approximately 12 g of powdered milk per 100 mL of liquid sample. The percentage of evaporated matter over time for a 100 mL crystallizer is illustrated in the Fig. 2.
@@ 2.3 Calcination
The powdered milk was transferred into a porcelain beaker and heated in a muffle for approximately 18 hours. Initially, calcination was performed at 550°C. However, organic matter removal was incomplete, yielding grayish ash (see Fig. 3). The colour of the ash serves as a reliable indicator of residual organic materials, with bright white signifying complete organic matter removal. Furthermore, calcined ash obtained at 550°C clogged the resin column or significantly reduced the flow rate.
@@ 2.4 Ion Exchange chromatography
The process begins with preconditioning the column using the same reagent as the sample solution. Since the calcined ash is diluted in 10 mL of 8M HNO3, the column is preconditioned by adding 5 mL of 8M HNO3.The diluted sample is then passed through the column. A series of rinses follows to separate strontium (stable or radioactive) from most other elements, based on their retention behaviour [6], which varies with nitric acid concentration. The rinsing sequence begins with 15 mL of 8M HNO3, followed by 5 mL of 3M HNO3 mixed with 0.05M oxalic acid to eliminate potential actinides. This step is crucial, as oxalic acid facilitates the formation of soluble actinide complexes, preventing their retention on the Sr Resin® column [7]. Finally, 7 mL of 8M HNO3 is used to complete the rinsing process.
@@ 2.5 Liquid Scintillation Spectroscopy
Liquid scintillation measurements were performed using a Packard TriCarb 2900TR spectrometer. The radioactive liquid eluted from the chromatography column was mixed with Ultima Gold AB scintillation liquid inside a 20 mL borosilicate glass vial. Samples were measured immediately after the chromatography process (t=0), twenty days later (t=20 d) to ensure that the 90Sr/90Y pair reached secular equilibrium, and finally after 30 days (t=30 d). This approach allowed for a clearer distinction of each radioisotope’s contribution to the final spectrum shape, thereby optimizing the final activity calculation (see Fig. 4).
@@ 2.6 Mass Spectrometry
The process consisted of two main steps. The first step involved calibrating the instrument using five aqueous solutions spiked with a stable strontium tracer at various concentrations. These samples were prepared and measured to generate a calibration curve for strontium concentrations ranging from 0.1 to 10 ppm (see Fig. 5).
@@
Moreover, ICP-MS enables the detection of trace amounts of other elements within the sample. This capability was used to assess whether other elements could interfere with the activity values obtained from liquid scintillation spectroscopy.
@@ 3 Results and Discussion
As part of an intercomparison measurement between several certified Swiss nuclear laboratories, a milk sample containing 90Sr at an unknown concentration was received, known to be below the Swiss release limit for this radionuclide (1 Bq/mL, as defined in the Swiss Radiological Protection Ordinance, ORaP [9]). To replicate a similar 90Sr concentration to that expected in the received samples, reference milk samples were prepared with a concentration of 0.45 Bq/mL. The protocol was then applied, achieving a high efficiency of 1002% as measured by ICP-MS, demonstrating the method’s reproducibility and robustness. These results confirm that the developed protocol retains 100% of the added 90Sr, enabling accurate quantification through liquid scintillation spectroscopy.
@@
To further validate the protocol’s accuracy, more LSC measurements were conducted at 20 and 30 days to track the increase in 90Y activity until secular equilibrium was reached. At this point, the absolute activity of all measured samples doubled and then stabilized, confirming the effective separation of strontium and yttrium during chromatography (see Fig. 6).
@@
The detection limit of 0.1 Bq was determined by measuring a series of 90Sr-spiked aqueous samples, ranging from 0.1 to 1 Bq (see Fig. 7), using the liquid scintillation spectrometer. These water-based standards were used to simulate the final eluates obtained after column separation, in order to assess the instrumental detection limit independently of the milk matrix. This approach allows for the calculation of the MDA under realistic laboratory conditions and highlights the role of detector sensitivity in determining overall method performance.
@@ 4 Conclusions and Outlook
This study presents a robust and efficient protocol for the quantification of 90Sr in milk, addressing critical food safety concerns related to potential radioactive contamination. The proposed method integrates freeze-drying, high-temperature calcination (at 700°C), ion exchange chromatography using Sr Resin®, and liquid scintillation spectroscopy. It achieves a high recovery efficiency of 1002%, as determined by ICP-MS measurements. Theoretical calculations indicate a Minimum Detectable Activity (MDA) of 0.33 Bq/L under optimal conditions. Additionally, the protocol was experimentally validated at a concentration of 0.45 Bq/mL, demonstrating its practical suitability for detecting low-level 90Sr contamination in milk, particularly during nuclear emergencies.
Beyond its performance metrics, the strength of this protocol lies in the reliability and reproducibility of its core steps, which include freeze-drying, calcination, and strontium-specific separation. These stages are essential for the elimination of organic matter and the selective extraction of strontium, and they can be applied to a broad range of radionuclides in food matrices. Unlike chemical digestion or precipitation techniques, this protocol relies mainly on physical processing steps and minimal reagent use, making it safer to implement, less operator-dependent, and more cost-effective. Furthermore, Sr Resin® columns can be reused multiple times without significant loss of performance, further reducing operational costs [12, 13, 14].
Comparative testing with other methods such as the Guérin and Dowex® 50W-X8 protocols highlighted the challenges of achieving high and consistent yields while ensuring safe handling conditions. The proposed method was developed with a focus on reproducibility, safety, and ease of use, especially when dealing with radioactive samples. While other methods may offer shorter processing times, they often require intensive manual handling or multiple centrifugation steps, which can limit scalability and introduce variability.
Although designed with nuclear emergency contexts in mind, this protocol is especially well suited for routine monitoring of large sample batches. Freeze-drying and calcination are easily scalable and compatible with high-throughput workflows, unlike centrifugation-based methods, which require repeated manual interventions to process larger volumes. Additionally, while calcination may appear time-consuming, the process involves primarily equipment time, not operation time, reducing workload and exposure.
Future work could adapt this protocol to other food matrices such as vegetables, grains, or processed products. Pairing it with more advanced detection systems such as ultra-low-level liquid scintillation counters would further improve sensitivity and extend its applicability to very low contamination levels. Overall, this method offers a practical, safe, and scalable framework for reliable 90Sr monitoring in food, contributing to improved preparedness and public health protection in radiological contexts.
@@ Acknowledgements
The authors would like to express their sincere gratitude to the inTECH institute of the School of Landscape, Architecture, and Engineering (HEPIA), of University of Applied Sciences and Arts Western Switzerland for their invaluable support throughout this research. The provision of equipment and laboratory time to our nuclear physics and nuclear chemistry laboratories was instrumental in the successful completion of this study.